# UniHENN: Designing More Versatile Homomorphic Encryption-based CNNs without im2col

### v1

@@ VIII Limitations
We propose an optimized CNN model inference mechanism based on HE, UniHENN. UniHENN uses batching to reduce inference time, but it has three limitations.
As deep learning has evolved, it has been used in many applications beyond image classification [48], such as detecting attacks on the web [45] or analyzing resource extraction [47]. These DNN models require a large number of operations, especially multiplication operations. To apply homomorphic encryption, multiplication operations must be used within a certain number of times depending on the parameter values used, or bootstrapping must be used. Due to the nature of deep learning models with a large number of multiplication operations, bootstrapping is required, but bootstrapping has the limitation of slowing down the computation speed significantly.
UniHENN does not support multi-core architecture or GPUs. As we tried to deploy our source codes, we chose to use SEAL-Python, an open-source homomorphic encryption framework. But most open-source homomorphic encryption frameworks, including SEAL-Python, are suitable for only CPU and single-thread architecture. Nowadays, some studies for optimizing HE operations in the GPU environment [6, 40] are conducted. The main idea of the study is to improve performance of the HE-operation using the parallelism design. So, to address these limitations, we plan to extend our architecture that support the multi-core environment and GPU environment in the software side with the previous works for future work.
Since UniHENN utilizes the batch system to reduce computational time. UniHENN will be efficient when analyzing sensitive data at scale or when clients infer multiple data at once. However, UniHENN may not be efficient when it is needed to infer single data because the design point of UniHENN is focused on batch operations. To overcome this limitation, we plan to improve the computational efficiency that the operations constituting the convolutional layer or FC layer operation will be distributed in the space where the batch operation was performed to increase efficiency.
@@ IX Related Work
Dowlin et al. [17] proposed CryptoNets, the first HE-based CNN inference scheme. It utilizes square activation instead of ReLU to reduce the total computation cost with a low number of multiplications. However, the input ciphertext size and weights of CryptoNets are very large because they encrypt each pixel of the input image and each element of the weights individually. Chillotti et al. [12] introduced an enhanced version of CryptoNets, applying low-degree polynomials for functions like ReLU, Softplus, and Swish and incorporating techniques like pruning and quantization for more efficient homomorphic operations. Ishiyama et al. [20] adopted an approximation of Google’s Swish activation function using two and four-degree polynomials, coupled with batch normalization, aiming to minimize the accuracy gap between traditional and polynomial-based activation functions. In 2021, Benaissa et al. presented TenSEAL [7], a prominent open-source tool that supports HE-based CNN inference. It utilizes the im2col method to efficiently compute the convolutional layer. However, it cannot support multiple convolutional layers and does not support batch operation. In this paper, we propose UniHENN to address these issues.
Aharoni et al. [4] presented an encryption data structure, the tile tensor, for efficiently calculating deep neural networks containing CNNs. It splits and encrypts the input data considering the model structure. However, the encryption of input data in this scheme depends on the model architecture; sharing the client’s ciphertext with a third party, such as a public cloud service, is challenging because the CNN model’s structure is usually undisclosed. Lee et al. [32] proposed a novel low-latency model optimization strategy using an enhanced HerPN block and masking convolution, significantly reducing multiplication operations in the convolution layer. However, their approach does not consider the design of FC layers. In this paper, we present how to implement all necessary layers for CNNs in HE efficiently.
@@ X Conclusion
In this paper, we introduce a privacy-preserving machine learning framework, UniHENN, addressing a significant challenge in cloud computing. UniHENN enables the construction of HE-enabled CNN models without relying on the traditional im2col operation, enhancing compatibility with diverse ML models and reinforcing privacy in cloud-based machine learning.
Our extensive evaluation of UniHENN across four public datasets using seven different CNNs, including six 2D CNNs and a 1D CNN, demonstrates its robustness and efficacy. Notably, UniHENN maintains accuracy levels comparable to unencrypted models, highlighting its practical applicability without performance compromise.
A key feature of UniHENN is its batch mechanism support, significantly accelerating processing times for multiple datasets concurrently, a breakthrough for cloud computing. Our findings indicate that a CNN constructed with UniHENN outperforms the model $M1$ (comprising one convolutional layer and two fully connected layers) by almost 3.921 times using the TenSEAL library, demonstrating a substantial improvement in processing speed, crucial for real-world applications requiring rapid data processing. We publicly release UniHENN’s source code at https://github.com/hm-choi/uni-henn to enhance its impact, promoting widespread use.
In the future, we aim to optimize UniHENN for multi-thread architectures, leveraging the additional space available in batch computations. This optimization will involve dividing the operations of each layer to enable simultaneous processing, thereby enhancing the efficiency of UniHENN. We aim to establish UniHENN as a practical tool for secure and efficient cloud computing in machine learning applications.

### v3

@@ VIII Limitations
We propose an optimized CNN model inference mechanism based on HE, UniHENN. While UniHENN uses batching to reduce inference time, it has three key limitations.
Firstly, complex deep learning models often require a large number of operations, particularly multiplications. In HE, multiplication operations must be limited based on parameter values, or bootstrapping must be used. Due to the extensive multiplication operations in deep learning models, bootstrapping is necessary, significantly slowing computation speed.
Secondly, UniHENN does not currently support multi-core architectures or GPUs. Our deployment uses SEAL-Python, an open-source HE framework suited for CPU and single-thread architectures. Most open-source HE frameworks, including SEAL-Python, are designed this way. Recent studies have optimized HE operations for GPUs [6, 42], focusing on parallelism to enhance performance. To address these limitations, we plan to extend our architecture to support multi-core and GPU environments. Our future work will involve optimizing key HE operations for parallel execution on GPUs, implementing efficient memory management techniques, and ensuring proper synchronization mechanisms to maximize computational performance.
Finally, UniHENN is designed to reduce computational time through batching, making it efficient for analyzing sensitive data at scale or when inferring multiple data points simultaneously. However, due to its focus on batch operations, UniHENN may be inefficient for inferring single data points. To overcome this, we plan to improve computational efficiency by distributing operations within the convolutional and FC layers in the space where batch operations are performed, thereby increasing overall efficiency.
@@ IX Related Work
IX-A Libraries of HE
Several prominent HE libraries are available, including SEAL [46], Lattigo [40], HElib [1], and OpenFHE [5]. Each offers unique features: SEAL [46] supports BFV and CKKS schemes with optimizations like ciphertext packing. We used SEAL-Python, a Python port of the C++ implementation, for our work. Lattigo [40] implements Ring-LWE-based HE primitives and Multiparty-HE protocols in Go, enabling cross-platform builds and optimized secure computation. HElib [1] supports BGV and CKKS schemes with bootstrapping, focusing on Smart-Vercauteren ciphertext packing and Gentry-Halevi-Smart optimization. It also provides an HE assembly language for low-level control. OpenFHE [5] offers bootstrapping and hardware acceleration using a standard Hardware Abstraction Layer (HAL).
IX-B Machine learning implementation with HE
To provide privacy-preserving ML services, many previous studies implement HE-friendly neural networks. They focus on two challenges: “How to build HE-friendly model architecture, especially, the activation layer” and “How to build an efficient HE-based CNN (or DNN) system.”
In the first case, there were efforts to replace activation functions with polynomial functions. CryptoNets [19] use a square function to replace the ReLU activation function. Chou et al. [13] proposed some techniques to replace activation functions such as ReLU, Softplus, and Swish with low-degree polynomials and use pruning and quantization for the efficiency of homomorphic operations. Ishiyama et al. [22] adopt Google’s Swish activation function with two- and four-degree polynomials with batch normalization to reduce the errors between Swish and approximated polynomials. However, replacing activation functions with low-degree polynomials reduces model accuracy and is only applicable to shallow models. To overcome these limitations, Park et al. [43] proposed the HerPN block, which can replace the batch normalization and ReLU block by utilizing the Hermite polynomial. Recently, Lee et al. [33] proposed PP-DNN, a low-latency model optimization solution focusing on convolution and approximate ReLU operations without bootstrapping, supporting CNN models like ResNet-34 with HE. In contrast, UniHENN offers comprehensive optimizations across all layer types, including convolutional, pooling, fully connected, and flatten layers, resulting in more efficient operations overall. Additionally, UniHENN introduces a batch system for multi-input inference and optimizes element sizes for batch operations, which are not addressed in PP-DNN.
In the second case, many studies [23, 43, 18] interact with clients to compute non-linear operations. The server sends the encrypted intermediate results to the client, and the client computes the non-linear functions with decrypted intermediate results. In this way, the model accuracy can be preserved, but the communication overhead is increased.
@@ X Conclusion
[…]